Two seniors, one from each government class are randomly selected to travel to Washington, DC. John is in a class of 18 students and Maureen is in a class of 20 students. Find the probability that both John and Maureen will be selected.
Answer by checkley77(12844)

1/18*1/20=1/360=0.002777=.28% ANS.